OpenStudy (anonymous):

I will fan and medal if you can help me with this!!! Come up with a new linear function that has a slope that falls in the range -1 < m < 0. Choose two different initial values. For this new linear function, what happens to the function’s values after many iterations? Are the function’s values getting close to a particular number in each case?

OpenStudy (anonymous):

@SolomonZelman Can you help me out please?

OpenStudy (anonymous):

@KJSaif please help me!

OpenStudy (solomonzelman):

Choose any number in between (and not including) -1 and 0.

OpenStudy (solomonzelman):

A linear function is in a form: \(\color{#000000}{ \displaystyle y=mx+b }\) Where \(\color{#000000}{ \displaystyle m }\) is the slope (or the rate of change, or the rate of growth), and \(\color{#000000}{ \displaystyle b }\) is the y-intercept (or some starting value.

OpenStudy (solomonzelman):

So, you can choose any \(\color{#000000}{ \displaystyle m }\) that is between 0 and -1, and whatever \(\color{#000000}{ \displaystyle b }\) that you feel like.